# Tokenization Falling Short: hTe Cusre of Tkoeniaztion

### v1

@@ E.1 Impact of Typographical Variations on Sequence Length
Figure 12 shows a strong positive correlation between token lengths before and after introducing typographical errors across different tasks (GSM8K and MMLU) and $n$-gram settings (2, 3, 5). We observe that token lengths after introducing errors are proportional to their original lengths. Besides, both GSM8k and MMLU tasks exhibit similar patterns, and the $n$-gram settings (2, 3, 5) do not significantly alter this relationship. Most Most interestingly, the slope for reorder errors is relatively larger than for noise errors, indicating that reorder errors tend to result in a slightly greater increase in token length compared to noise errors.
@@ E.2 Compositional Challenges in Token Embeddings
In Figure LABEL:fig:vec_assignment, we observe the word "assignment" and its subword components "assign" and "ment." The cosine similarity between the composite embedding "assign + ment" and the original word "assignment" is relatively low at 0.51, with a significant angular difference of 59.24 degrees. This substantial disparity indicates that the learned token embeddings fail to capture the surface form composition accurately. The model does not recognize that "assign" combined with "ment" should semantically align closely with "assignment."
Similarly, Figure LABEL:fig:vec_import shows the word “import” and its subword components “im” and “port”. Here, although the cosine similarity between “im + port” and “import” is higher at 0.93, the angular difference is still notable at 21.17 degrees. This suggests that while the model captures some compositional aspects, it still struggles to fully integrate subword information to match the original word’s embedding perfectly.

### v2

@@ E.1 Impact of Typographical Variations on Sequence Length
Figure 12 shows a strong positive correlation between token lengths before and after introducing typographical errors across different tasks (GSM8K and MMLU) and $n$-gram settings (2, 3, 5). We observe that token lengths after introducing errors are proportional to their original lengths. Besides, both GSM8k and MMLU tasks exhibit similar patterns, and the $n$-gram settings (2, 3, 5) do not significantly alter this relationship. Most interestingly, the slope for reorder errors is relatively larger than for noise errors, indicating that reorder errors tend to result in a slightly greater increase in token length compared to noise errors.
@@ E.2 Compositional Challenges in Token Embeddings
In Figure LABEL:fig:vec_assignment, we observe the word “assignment” and its subword components “assign” and “ment”. The cosine similarity between the composite embedding “assign + ment” and the original word "assignment" is relatively low at 0.21, with a significant angular difference of 78.16 degrees. This substantial disparity indicates that the learned token embeddings fail to capture the surface form composition accurately. The model does not recognize that “assign” combined with “ment” should semantically align closely with “assignment”.
Similarly, Figure LABEL:fig:vec_import shows the word “import” and its subword components “im” and “port”. Here, the cosine similarity between “im + port” and “import” is 0.13, and the angular difference is still notable at 82.47 degrees. This suggests that while the model captures some compositional aspects, it still struggles to fully integrate subword information to match the original word’s embedding perfectly.